Summary for OWL, Friday, August 20, 1999

Angela Bellavance (bellavan@rice.edu)
Fri, 20 Aug 1999 08:19:51 -0500 (CDT)

I'm not sure this actually got sent out from the previous shift, so here
it is again...
*********************************************************************

Shift Summary for Friday OWL shift, Aug 20, 1999

On Shift: R.Kessler, A.Erwin

Hours of beam available at all intensities: 6.9 hrs
Hours of beam available at nominal intensity: 6.9
Hours of E832 data (special runs not included): 6.6 hrs
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:
10 minutes from Pipeline problems at beginning of new run
10 minutes from csi patch going blank [pipeline ??]

========================================================================
Runs ended this shift:
Run 14277 BEAM_832 in progress

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: NM2SR is ramping to about 140 Amps below the set value.
Rick Ford was sent an email. Running was too smooth
to risk calling MCR.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:00 Begin new run shortly after midnight. Previous
shiftees graciously stayed overtime and fixed
the pipeline mess that started at the beginning
of a new run. Rebooted KTPIPE3 with Ram's help &
cycled power.

14277 132 03:03 pipeline problem. Reinit pipeline and Kfrend.
136 03:06 restart trigger

3:30 Pre-print of e'/e lattice calculation sent to ktev
printer. Shift crew decides to check computation
of <2pi|Q6|K> operator using ktev4. However,
too much DAQDEAD time forces us to kill the job.

237 5:22 Segment of CsI disappears on event display.
Reinit pipeline.

242 Good spill again.

6:49 Angela arrives early.
7:03 Taku Arrives LATE.
(AMB: Actually it was only 7:01!)

========================================================================
Follow up Notes on Previous Shift: none

========================================================================
Items to Follow up on Next Shift:
o Watch low NM2SR currents
o read lattice theory paper on e'/e.
It's a theory paper with 14 authors.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke